OpenStudy (anonymous):

Find the slope of the line through (-5, -2) and (8, 10).

OpenStudy (anonymous):

M=y2-y1/x2-x1 Use this formula and get the answer

OpenStudy (texaschic101):

we will be using the slope formula : slope = (y2 - y1) / (x2 - x1) we have 2 sets of points.. 1st set (-5,-2).....so we will call x1 = -5 and y1 = -2 2nd set (8,10)...so we will call x2 = 8 and y2 = 10 now we sub into the formula slope = (10 -(- 2) / (8 - (-5) slope = (10 + 2) / (8 + 5) slope = 12/13 any questions ?

OpenStudy (anonymous):

@satellite73. Now would u not have any question for above this 99 smart score person?